OpenStudy (anonymous):

combine and simplify (1/3)(cube root of 81)

OpenStudy (anonymous):

\[\sqrt[3]{81}=\sqrt[3]{9\times9}=\sqrt[3]{3\times3\times3\times3}\]

OpenStudy (anonymous):

So now we can pull a three out from under the radical since it is a cube root and we have at least 3 3s under the radical

OpenStudy (anonymous):

\[\sqrt[3]{3\times3\times3\times3}=3\sqrt[3]{3}\]

OpenStudy (anonymous):

Now we multiply this by 1/3\[1/3\times3\sqrt[3]{3}=\sqrt[3]{3}\]

OpenStudy (anonymous):

Hope this helps :)